# Tackling GenAI Copyright Issues: Originality Estimation and Generecization

### v5

@@ Abstract
The rapid progress of generative AI technology has sparked significant copyright concerns, leading to numerous lawsuits filed against AI developers. While various techniques for mitigating copyright issues have been studied, significant risks remain. Here, we propose a genericization method that modifies the outputs of a generative model to make them more generic and less likely to infringe copyright. To achieve this, we introduce a metric for quantifying the level of originality of data in a manner that is consistent with the legal framework. This metric can be estimated by drawing samples from a generative model, which is then used for the genericization process. As a practical implementation, we introduce PREGen, which combines our genericization method with an existing mitigation technique. Experiments demonstrate that our genericization method successfully modifies the output of a text-to-image generative model so that it produces more generic, copyright-compliant images. Compared to the existing method, PREGen reduces the likelihood of generating copyrighted characters by more than half when the names of copyrighted characters are used as the prompt, dramatically improving the performance. Additionally, while generative models can produce copyrighted characters even when their names are not directly mentioned in the prompt, PREGen almost entirely prevents the generation of such characters in these cases.
@@ 1 Introduction
Generative models have exhibited performance that rivals human capabilities in creative tasks such as those involving image synthesis and language processing [rombach2022high, bubeck2023sparks]. However, this progress has also raised concerns about copyright protection, leading to numerous lawsuits filed by creators against AI developers [samuelson2023]. Copyright law aims to protect the rights of creators over their works, encouraging the creation of new works while maintaining the balance between the interests of creators and the public [litman1990]. Concerns about copyright issues with generative models arise primarily because these models are trained on massive datasets that often contain copyrighted works and are capable of producing outputs closely resembling their training data, potentially resulting in violation of exclusive rights of copyright owners [sag2023].
@@
In this paper, we propose a method for quantifying the level of originality and modifying the outputs of generative models to those that have lower originality values. These modified outputs are more generic and less likely to infringe copyright. As a proof of concept, we apply our genericization method to a text-to-image generative model to demonstrate that it successfully concentrates the outputs of the generative model around more generic, copyright-compliant ones while suppressing outputs that are highly similar to copyrighted images.
@@
Focusing on the generation of copyrighted characters, which poses particular copyright challenges [sag2023], we demonstrate that PREGen significantly improves the performance of prompt rewriting accompanied by a nagative prompt (we refer to this as the standard method when there is no ambiguity). In particular, experiments using the COPYCAT benchmark [he2024fantasticcopyrightedbeastsnot] show that when copyrighted characters’ names are used as the prompt, PREGen reduces the likelihood of text-to-image generative models generating images of copyrighted characters by more than half compared to the standard method. Furthermore, although generative models can inadvertently generate copyrighted characters even with seemingly harmless prompts [he2024fantasticcopyrightedbeastsnot, lee2024talkin], PREGen almost completely eliminates the generation of copyrighted characters in such scenarios.
2 Quantifying originality
@@
AI developers and users might infringe copyright if generative AI produces outputs that resemble existing works. Specifically, outputs that are substantially similar to original works can potentially infringe copyright in them [crs2023, lee2024talkin]. Consequently, infringement is found only when the significance of similarity reaches some threshold, involving both qunatitative and qualitative aspects, that it is considered substantial [balganesh2012]. Copying must be found on protectable elements of the original work [samuelson2015], and common or standard expressions are usually deemed unprotectible via application of various legal doctrines. For example, scènes à faire denies copyright protection to works or parts of works that are highly typical or commonly found within the specific genre to which the work belongs [lund2009, hacohen2024]. The substantial similarity test operates in such a way that the higher the level of originality a work possesses–in the sense that it involves more creative and expressive qualities [^1] The meaning of originality here is distinct from the basic requirement under copyright law that a work be independently created and display a minimal level of creativity, which qualifies it as an “original work of authorship” eligible for protection.–the thicker its copyright protection. This means that broader range of copying may be considered infringment. Conversely, works with less originality receive thinner protection, meaning fewer types and forms of copying are likely to be considered infringing [balganesh2012, vermont2012].
@@
Aside from the likelihood and degree of being distinct from other works, as captured in our metric, a common-sense notion of originality might encompass such factors as novelty, ingenuity and artistic merit. However, it is generally understood that the copyright law does not directly concern these qualities of the work [gorman2001, vermont2012]. Thus, in principle, it is appropriate not to take them into account when determining the strength of copyright protection beyond the extent that is already captured by our metric.
@@
The distance metric $d$ involved in Equation 1 could be some quantity that is inversely related to the similarity between the works. Assessment of similarity in the context of copyright involves an intuitive judgment [litman1990], and the myriads of standards and methods devised by courts to analyze the substantial similarity requirement could partly be understood as tools to provide explanations to ad hoc intuitive judgments [hickey2016]. Such intuition could be captured by deep learning-based methods, which have been found to align well with human perception [zhang2018, reimers2019].
In our experiments, we use the cosine distance in terms of CLIP [radford2021] and DINOv2 [oquab2024dinov2] feature embeddings. In equations, cosine distance is defined as $d(a,b)=1-s(a,b)$ where $s$ is the cosine similarity defined as $s(a,b)=\frac{v_{a}\cdot v_{b}}{\rVert v_{a}\rVert\rVert v_{b}\rVert}$ with $v_{a},v_{b}$ being the feature embeddings of $a$ and $b$. Our framework allows for the flexibility of choosing the appropriate distance metric depending on the application.
@@ 3 Genericization
The resulting output is expected to be generic, as opposed to highly unique or original, relative to the prompt. As long as the prompt describes an object that is sufficiently generic, the final output will likely be generic. By combining this genericization method with prompt rewriting, we can genericize the outputs even when the prompt directly refers to or closely describes copyrighted content (see Section 5).
@@
The underlying philosophy of our method is to avoid infringement by generalizing outputs, ensuring they are not overly specific to any protected work. As a consequence of the varying level of copyright protection as embodied in the substantial similarity test, generic expressions–those that use common or widely shared ideas and themes with limited originality–are less likely to infringe on other copyrighted works. This approach is effective with copyrighted characters too, whose unique features make it easier to meet the threshold for substantial similarity [sag2023], while it has broader applicability to other protected works.
4 Proof of concept
Procedures
We use SDXL [podell2023sdxl], a text-to-image generative model, to obtain originality estimates (Equation 2) of two images of copyrighted characters, Mario from Super Mario Bros and Winnie the Pooh. We used five different prompts (Table 1) for each image as the conditioning $x$, designed to describe the characters at different levels of abstractness or specificity. The specific prompts are designed to capture the essence of the visual aspects of the copyrighted characters, and other prompts were formed by excerpting parts of the specific prompts.
We also test our genericization method by generating multiple samples and selecting the genericized outputs based on Equation 3. This process was applied to the outputs of SDXL created using the specific prompts listed in Table 1.
Mario Pooh
Abstract “a mustachioed character” “a bear”
Images of copyrighted characters as well as the sample images produced by the generative model using each prompt are displayed in Figure 1. More specific prompts tend to produce images that share more visual elements with the copyrighted characters. See Appendix F for further details of the experimental settings.
Figure 1: Images of copyrighted characters and characters generated by the generative model. The leftmost column shows the images of copyrighted characters for which originality values were computed. The subsequent columns display sample images generated by the generative model, starting from those generated using the abstract prompts on the left and moving towards those generated with increasingly specific prompts towards the right.
Results
The results of originality estimation and genericization using CLIP are presented. The results with DINOv2 are similar and are shown in Appendix A.
Originality estimation
Our metric successfully assigns higher originality values when the generative model is given more abstract prompts and lower originality values for more specific prompts. This reflects the fact that typical generation with more specific prompts, which more closely describe the copyrighted images, is more likely to be highly similar to copyrighted images. In addition, originality estimates for images of copyrighted characters tend to be higher than typical output from the generative model, especially when more abstract prompts are used (Figure 2). However, our originality estimation method can fail when the distribution learned by the generative model is highly distorted (see Appendix C for a case of failure). PREGen discussed in Section 5 addresses this issue by introducing diversity in the prompts used for internally generated samples.
Genericization
5 Practical implementation and perfomance
PREGen
We propose an algorithm to mitigate copyright risks of generative models (Algorithm 1), PREGen (Prompt Rwriting-Enhanced Genericization), [^3] The code is available at: https://github.com/hirochok/PREGen. that combines our genericization method involving originality quantification with the standard prompt rewriting method [he2024fantasticcopyrightedbeastsnot].
@@ 5 Practical implementation and perfomance
The algorithm first, as in the standard prompt rewriting method, modifies the input prompt to a clean prompt, which is a textual description that does not refer to any copyrighted characters or elements using a large language model (LLM). The standard method would use the clean prompt, together with an appropriate negative prompt, to produce content. Our algorithm, instead of directly using the clean prompt, further generates multiple variations of input prompts by rewriting the clean prompt. Then, each generated input prompt is fed into the generative model with the negative prompt to produce samples internally. The algorithm outputs the generation that has the lowest originality estimate among the internally produced samples.
@@
2:Initialize an empty set $S_{L}$ for storing prompts to be generated by $L$
@@
Performance
@@
In addition to DETECT and CONS, to aid our analysis, we define p-CONS, a new metric that directly measures the alignment of the generation with the input prompt provided by the user. p-CONS can be considered as a more fine-grained measure of prompt consistency than the standard CONS. The exact definition of p-CONS and the details of the anlysis involving p-CONS can be found in Appendix D.
@@
“Mario” “A short, plump man sporting a cap and overalls, both in a vibrant shade of blue. He wears a red shirt and gloves, and is always seen with a thick, bushy moustache. His rosy cheeks, bright blue eyes, and brown loafers complete his look.”
@@ Results
Table 4: Performance in the indirect anchoring scenario. Similar to the direct anchoring case, PREGen mostly outperforms the standard method. Only the CONS score for SDXL is lower than the standard method, but by a small margin. The DETECT scores drop to 0, no detection of copyrighted characters at all, for Playground v2.5 and SDXL.
@@
Figure 4: Images generated in the direct anchoring scenario. When the user’s prompt is the name of the copyrighted character itself, the generative models can generate images that closely resemble the copyrighted character. The standard prompt rewriting method (with negative prompting) modifies the outputs to those that are less similar to copyrighted characters. PREGen tends to successfully exclude elements that are highly unique to copyrighted characters even when the standard method fails to do so.
@@ 6 Related works
Researchers have recently explored the application of computational methods to copyright legal issues. For example, a framework for testing substantial similarity based on Kolmogorov-Levin complexity has been proposed [scheffler2022]. The potential of machine learning to bring objectivity to the substantial similarity test has also been recognized in the legal literature [tu2021]. While a deep-learning-based method for measuring creativity has been suggested [franceschelli2022], such methods can only be effective for our purposes if they align with the framework of copyright law. More recently, the ability of generative models to capture the distribution of expressions in a massive dataset has been considered promising for measuring originality [hacohen2024, hacohen2024similarities, haviv2024imageworththousandwords]. In this context, a method involving text inversion and reconstructing the original image from the resulting text tokens has been discussed as a means to measure the originality of images [hacohen2024, haviv2024imageworththousandwords]. This paper contributes to the literature by introducing a novel metric for quantifying the level of originality, which is consistent with the legal framework and can be practically estimated using generative models.
@@ 7 Discussion
We introduced a method to genericize the output of generative models, thereby reducing the risk of copyright infringement. Our method leverages the principle that the level of originality of works determines the strength of their copyright protection, as well as the capability of generative models to learn the distribution of training data. Experiments demonstrate that our genericization method effectively steers the outputs away from those highly similar to unique copyrighted data and produces more generic, copyright-compliant ones.
We further proposed PREGen, a practical algorithm that combines our genericization method with prompt rewriting. By evaluating the performance of PREGen using the COPYCAT suite, we have shown that PREGen significantly enhances the performance of the standard prompt rewriting method.
Our work is not without limitations. While the general framework for originality estimation and genericization is broadly applicable, our experiments have primarily focused on text-to-image generative models and distance metrics suited to them. Future research should investigate generative models and distance metrics for other types of generated content, such as text and video.
@@ Appendix E The effect of increasing parameter $b$
As we increase the value of parameter $b$ of PREGen, the performance tends to improve. In particular, CONS and p-CONS increase with no discernible negative impact on DETECT, both in the direct anchoring and indirect anchoring scenarios (Figure 10). Although PREGen’s CONS scores are lower than the standard prompt rewriting method at $b=1$, they exceed or match the performance of the standard method at $b=2$. It is suggested that, although the introduction of diversity in prompts via the prompt rewriting step in PREGen may deviate generations away from the user’s original propt, such deviations get averaged out as the number of batches increases.

### v6

@@ Abstract
The rapid progress of generative AI technology has sparked significant copyright concerns, leading to numerous lawsuits filed against AI developers. Notably, generative AI’s capacity for generating copyrighted characters has been well documented in the literature, and while various techniques for mitigating copyright issues have been studied, significant risks remain. Here, we propose a genericization method that modifies the outputs of a generative model to make them more generic and less likely to imitate distinctive features of copyrighted materials. To achieve this, we introduce a metric for quantifying the level of originality of data, estimated by drawing samples from a generative model, and applied in the genericization process. As a practical implementation, we introduce PREGen (Prompt Rewriting-Enhanced Genericization), which combines our genericization method with an existing mitigation technique. Compared to the existing method, PREGen reduces the likelihood of generating copyrighted characters by more than half when the names of copyrighted characters are used as the prompt. Additionally, while generative models can produce copyrighted characters even when their names are not directly mentioned in the prompt, PREGen almost entirely prevents the generation of such characters in these cases.
@@ 1 Introduction
Generative models have exhibited performance that rivals human capabilities in creative tasks such as those involving image synthesis and language processing [rombach2022high, bubeck2023sparks]. However, this progress has also raised concerns about copyright protection, leading to numerous lawsuits filed by creators against AI developers [samuelson2023]. Copyright law aims to protect the rights of creators over their works, encouraging the creation of new works while maintaining the balance between the interests of creators and the public [litman1990]. Concerns about copyright issues with generative models arise primarily because these models are trained on massive datasets that often contain copyrighted works and are capable of producing outputs closely resembling their training data, potentially resulting in violation of exclusive rights of copyright owners [sag2023]. This issue is especially significant in the generation of copyrighted characters, as models can easily create images sufficiently similar to these characters to infringe copyright, and they do so even when the characters’ names are not explicitly mentioned in prompts [sag2023, lee2024talkin, he2024fantasticcopyrightedbeastsnot].
@@
We demonstrate that PREGen significantly improves the performance of prompt rewriting accompanied by a negative prompt (we refer to this as the standard method when there is no ambiguity) in reducing the likelihood of generating images of copyrighted characters. In particular, experiments using the COPYCAT benchmark [he2024fantasticcopyrightedbeastsnot] show that when copyrighted characters’ names are used as the prompt, PREGen reduces the likelihood of text-to-image generative models generating copyrighted characters by more than half compared to the standard method. Furthermore, when the prompt does not directly reference copyrighted characters, PREGen almost completely eliminates the generation of copyrighted characters.
2 Originality estimation
@@
AI developers and users might infringe copyright if generative AI produces outputs that resemble existing works. Specifically, outputs that are substantially similar to original works can potentially infringe copyright in them [crs2023, lee2024talkin]. Consequently, infringement is found only when the significance of similarity reaches some threshold, involving both qunatitative and qualitative aspects, that it is considered substantial [balganesh2012]. Copying must be found on protectable elements of the original work [samuelson2013], and common or standard expressions are usually deemed unprotectible via application of various legal doctrines. For example, scènes à faire denies copyright protection to works or parts of works that are highly typical or commonly found within the specific genre to which the work belongs [lund2009, hacohen2024]. The substantial similarity test operates in such a way that the higher the level of originality a work possesses–in the sense that it involves more creative and expressive qualities [^1] The meaning of originality here is distinct from the basic requirement under copyright law that a work be independently created and display a minimal level of creativity, which qualifies it as an “original work of authorship” eligible for protection.–the thicker its copyright protection. This means that broader range of copying may be considered infringment. Conversely, works with less originality receive thinner protection, meaning fewer types and forms of copying are likely to be considered infringing [balganesh2012, vermont2012, hacohen2024].
@@
The metric embodies the principle that, as a work becomes more widespread and commonly used, its originality is perceived to decrease under copyright law. By diminishing protection for basic expressive elements–even if they were innovative and highly original when introduced–that have become frequently replicated and deeply embedded in culture, copyright law strikes a balance between encouraging originality and promoting broad access to foundational creative components [lund2009, hacohen2024]. Aside from the likelihood and degree of being distinct from other works, as captured in our metric, a common-sense notion of originality might encompass such factors as novelty, ingenuity and artistic merit. However, it is generally understood that the copyright law does not directly concern these qualities of the work [gorman2001, vermont2012]. Thus, in principle, it is appropriate not to take them into account when determining the strength of copyright protection beyond the extent that is already captured by our metric.
@@
The distance metric $d$ involved in Equation 1 could be some quantity that is inversely related to the similarity between the works. For image generation, possible choices of distance metrics include using between feature embeddings obtained from CLIP [radford2021] and DINOv2 [oquab2024dinov2], which achieve state-of-the-art performance in various tasks, such as image classification and representation learning. In equations, cosine distance is defined as $d(a,b)=1-s(a,b)$ where $s$ is the cosine similarity defined as $s(a,b)=\frac{v_{a}\cdot v_{b}}{\rVert v_{a}\rVert\rVert v_{b}\rVert}$ with $v_{a},v_{b}$ being the feature embeddings of $a$ and $b$. Our framework allows for the flexibility of choosing the appropriate distance metric depending on the application.
There are various forms of tests employed in case law for assessing substantial similarity, including the dissection approach, which breaks down works into specific expressive elements for detailed comparison; the total concept and feel approach, which evaluates the overall impression and combined elements of the works as a whole; and filtration methods, which first exclude unprotectable aspects before analyzing substantial similarities [samuelson2013]. Developing an automated method for these tests is challenging, and choices for the distance metric are made in part based on practicality.
@@ 3 Genericization
The resulting output is expected to be generic, as opposed to highly unique or original, relative to the prompt. The underlying philosophy of our method is to avoid infringement by genericizing outputs, ensuring they are not overly specific to any protected work. As a consequence of the varying level of copyright protection as embodied in the substantial similarity test, generic expressions–those that use common or widely shared ideas and themes with limited originality–are less likely to infringe on other copyrighted works.
4 PREGen
Our methods for originality estimation and genericization have limitations that render them inoperable when applied as-is. For the genericization method to work effectively, the input prompt must be sufficiently generic to yield non-infringing output and, for example, should avoid specific names of copyrighted characters. Additionally, even with a generic prompt, our originality estimation method can fail when the distribution learned by the generative model is highly distorted (see Appendix C for a case of failure). In such instances, the generated outputs may be predominantly infringing, and selecting the most “generic" outputs among them can even prove counterproductive. In light of these issues, we propose a practical algorithm to mitigate copyright risks of generative models (Algorithm 1), PREGen (Prompt Rewriting-Enhanced Genericization), [^3] The code is available at: https://github.com/hirochok/PREGen. which combines our genericization method based on originality quantification with the standard prompt rewriting method [he2024fantasticcopyrightedbeastsnot].
@@ 4 PREGen
The algorithm first, as in the standard prompt rewriting method, modifies the input prompt to a clean prompt, which is a textual description that does not refer to any copyrighted characters or hint at protectable elements using a large language model (LLM). The standard method would use the clean prompt, together with an appropriate negative prompt, to produce content. Our algorithm, instead of directly using the clean prompt, further generates multiple variations of input prompts by rewriting the clean prompt. Then, each generated input prompt is fed into the generative model with the negative prompt to produce samples internally. Finally, the algorithm outputs the generation that has the lowest originality estimate among the internally produced samples. The effectiveness of prompt rewriting is enhanced by further introducing a negative prompt that instructs the model not to generate specific content [he2024fantasticcopyrightedbeastsnot], an element that is also incorporated into PREGen.
@@
2:Initialize an empty set $\mathcal{S}_{L}$ for storing prompts to be generated by $L$
@@
5 Experiments
We first evaluate our originality estimation and genericization methods as-is, without integrating them into PREGen. Although this standalone approach is not practical for actual implementation for the reasons mentioned in 4, this experiment serves as a proof of concept to verify that the methods function as intended. We then test the performance of PREGen against an existing benchmark to test its practicality.
Originality quantification and genericization
Procedures
We use SDXL [podell2023sdxl], a text-to-image generative model, to obtain originality estimates (Equation 2) of two images of copyrighted characters, Mario from Super Mario Bros and Winnie the Pooh. We used five different prompts (Table 1) for each image as the conditioning $x$, designed to describe the characters at different levels of abstractness or specificity. The specific prompts are designed to capture the essence of the visual aspects of the copyrighted characters, and other prompts were formed by excerpting parts of the specific prompts. Images of copyrighted characters as well as the sample images produced by the generative model using each prompt are displayed in Figure 1. We also test our genericization method by generating multiple samples and selecting the genericized outputs based on Equation 3. This process was applied to the outputs of SDXL created using the specific prompts listed in Table 1.
The results of experiments using CLIP as the distance metric are presented here, while those using DINOv2 are similar and included in Appendix A. For further details of the experimental settings, see Appendix F.
Mario Pooh
Abstract “a mustachioed character” “a bear”
Figure 1: Images of copyrighted characters and characters generated by the generative model. The leftmost column shows the images of copyrighted characters for which originality values were computed. The subsequent columns display sample images generated by the generative model, starting from those generated using the abstract prompts on the left and moving towards those generated with increasingly specific prompts towards the right. More specific prompts tend to produce images that share more visual elements with the copyrighted characters.
Results for originality estimation
Our metric successfully assigns higher originality values when the generative model is given more abstract prompts and lower originality values for more specific prompts. This reflects the fact that typical generation with more specific prompts, which more closely describe the copyrighted images, is more likely to be highly similar to copyrighted images. In addition, originality estimates for images of copyrighted characters tend to be higher than typical output from the generative model, especially when more abstract prompts are used (Figure 2).
Results for genericization
Performance of PREGen
@@ 5 Experiments
In addition to DETECT and CONS, to aid our analysis, we define p-CONS, a new metric that directly measures the alignment of the generation with the input prompt provided by the user. p-CONS can be considered as a more fine-grained measure of prompt consistency than the standard CONS. The exact definition of p-CONS and the details of the analysis involving p-CONS can be found in Appendix D.
@@
“Mario” “A short, plump man sporting a cap and overalls, both in a vibrant shade of blue. He wears a red shirt and gloves, and is always seen with a thick, bushy mustache. His rosy cheeks, bright blue eyes, and brown loafers complete his look.”
@@ Results
Table 4: Performance in the indirect anchoring scenario. Similar to the direct anchoring case, PREGen mostly outperforms the standard method. Only the CONS score for SDXL is lower than the standard method, but by a small margin. The DETECT scores drop to zero, no detection of copyrighted characters at all, for Playground v2.5 and SDXL.
@@
Figure 4: Images generated in the direct anchoring scenario. When the user’s prompt is the name of the copyrighted character itself, the generative models can generate images that closely resemble the copyrighted character. PREGen tends to successfully exclude elements that are highly unique to copyrighted characters even when the standard method fails to do so.
@@ 6 Related works
Researchers have recently explored the application of computational methods to copyright legal issues. For example, a framework for testing substantial similarity based on Kolmogorov-Levin complexity has been proposed [scheffler2022]. The potential of machine learning to bring objectivity to the substantial similarity test has also been recognized in the legal literature [tu2021]. While a deep-learning-based method for measuring creativity has been suggested [franceschelli2022], such methods can only be effective for our purposes if they align with the framework of copyright law. As a concept for ensuring copyright safety in generative models, the near access-free condition has been proposed [vyas2023], although it has faced some criticisms [lee2024talkin, chibaokabe2024probabilistic]. More recently, the ability of generative models to capture the distribution of expressions in a massive dataset has been considered promising for measuring originality [hacohen2024, hacohen2024similarities, haviv2024imageworththousandwords]. In this context, a method involving text inversion and reconstructing the original image from the resulting text tokens has been discussed as a means to measure the originality of images [hacohen2024, haviv2024imageworththousandwords]. This paper contributes to the literature by introducing a novel metric for quantifying the level of originality, which is consistent with the legal framework and can be practically estimated using generative models.
@@ 7 Discussion
We introduced a method to genericize the output of generative models, thereby reducing the risk of copyright infringement. We further proposed PREGen, a practical algorithm for mitigating copyright risks, which combines our genericization method with prompt rewriting. Our method leverages the principle that the level of originality of works determines the strength of their copyright protection, as well as the capability of generative models to learn the distribution of training data.
By evaluating the performance of PREGen using the COPYCAT suite, we have shown that PREGen significantly enhances the performance of the standard prompt rewriting method. However, this improvement comes with a trade-off: PREGen requires additional computation to generate multiple samples, most of which are ultimately discarded, along with rewritten prompts. Additionally, consistency with the fine-grained details of the original prompt may be reduced.
Our work has certain limitations in its scope. While the general framework for originality estimation and genericization is broadly applicable, we have focused on the generation of copyrighted characters using text-to-image generative models. Future research can test our method on the generation of other types of materials and the use of different generative models, such as those for text and video, and investigate appropriate distance metrics and their effectiveness.
@@ Appendix E The effect of increasing parameter $b$
As we increase the value of parameter $b$ of PREGen, the performance tends to improve. In particular, CONS and p-CONS increase with no discernible negative impact on DETECT, both in the direct anchoring and indirect anchoring scenarios (Figure 10). Although PREGen’s CONS scores are lower than the standard prompt rewriting method at $b=1$, they exceed or match the performance of the standard method at $b=2$. It is suggested that, although the introduction of diversity in prompts via the prompt rewriting step in PREGen may deviate generations away from the user’s original prompt, such deviations get averaged out as the number of batches increases.